u Rydroxide is required to neutralize 43.00 mL of 0.400 M Answer #4 5. Analysis ofa solid acid shows its composition to be 2.22% hydrogen, 26.67% carbon, and 71.1190 oxygen. A 1.17gram sample is titrated, to neutralization, with 20 mL of a 0.65 M. Determine the molecular formula of the acid. Answer #5
Answer #1

1.17 g of sample of acid will contain :

hydrogen : (2.22 x 1.17) / 100 = 0.025974 g

Number of moles = 0.025974 / 1.007 = 0.02579 moles

carbon : (26.67 x 1.17) / 100 =0.312039 g

Number of moles = 0.312039 / 12.01 = 0.02598 moles

oxygen = ( 71.11 x 1.17) / 100 =0.831987 g

number of moles = 0.831987 / 32 = 0.025999 moles

Divide these number of moles by a common value , to get the ratio as 1:1:1 ,

So the empirical formula of the compound becomes = CHO

empirical mass = 12.01 + 1.007 + 32= 45.017 g

Number of moles of base = 0.65 x 0.020 = 0.013 moles

Since equimolar concentrations of both acid and base are used , so the number of moles of acid will also be 0.013 moles

So 0.013 moles of acid weighs 1.17 g

So molar mass will be : 1.17 / 0.013 = 90 g/mol

Divide molar mass by empirical mass to get the factor as 2

Now multiply the empirical formula by 2 , to get the molecular formula of the acid as : C2H2O2